# Don’t Forget Its Variance! The Minimum Path Variance Principle for Accurate and Stable Score-Based Models

### v2

@@ 6 Conclusions and Future Works
This paper addresses the key paradox in score-based density ratio estimation (DRE) that estimator performance strongly depends on the chosen path schedule while being theoretically path-invariant. Our analysis reveals that this discrepancy arises from an overlooked path variance term in training objectives. Based on this observation, we established the Minimum Path Variance (MinPV) principle, arguing that this term should be explicitly minimized rather than ignored. We then proposed MinPV, a framework that derives a closed-form expression for path variance and directly optimizes it. By parameterizing the path with a Kumaraswamy Mixture (KMM) Model, MinPV learns a data-adaptive low-variance path. This eliminates the need for heuristic path selection and ensures closer alignment with the ideal score-matching objective. Experiments demonstrate that MinPV achieves SOTA performance in estimation error and stability across challenging DRE benchmarks.
@@
The MinPV principle is general and could benefit other score-based generative models by learning optimal noise schedules to improve sample quality and stability. Beyond accuracy, future research could focus on computational efficiency, for example by leveraging optimal transport–based path formulations or adopting lighter parameterizations to further accelerate training. We believe that explicitly optimizing the geometry of interpolation paths will continue to afford significant advances in density ratio estimation and related score-based methods.
@@ A.1 Assumptions
A.5 Control of the Lipschitz Constant $L$ via Minimum Path Principle
@@ B.1 Formulations and Variances of Common Path Schedules
This path is inspired by the variance-preserving (VP) SDE schedule (Ho et al., 2020; Song et al., 2020a), satisfying the spherical constraint $\alpha(t)^{2}+\beta(t)^{2}=1$. We use the standard diffusion constants $\beta_{0}=0.1$ and $\beta_{1}=20$.
@@ B.3 Experimental Settings for the Kumaraswamy Distribution
PyTorch implementation of MinPV path with KMM parameterization.
We also provide the PyTorch implementation of the proposed MinPV in this paper, as shown in Algorithm 3.
Algorithm 3 PyTorch implementation of MinPV path with KMM parameterization.
1class MinPVPath($n$.Module):
@@
3 An implementation for computing the MinPV path schedules
@@ B.4 Experimental Settings and Results for $f$-divergence Estimation
    Edge-Singular Gaussian. This task assesses model stability in edge-case correlation scenarios. We consider two jointly Gaussian random variables, $({\mathbf{x}},{\mathbf{y}})\sim\mathcal{N}(\mathbf{0},[[1,\rho],[\rho,1]])$. The mutual information, given by $I({\mathbf{x}};{\mathbf{y}})=-0.5\log(1-\rho^{2})$, becomes numerically challenging at the extremes of the correlation coefficient $\rho$. As $\rho\to 1^{-}$, the distributions become nearly degenerate and collinear, leading to numerical divergence in many estimators. Conversely, as $\rho\to 0^{+}$, the dependency signal becomes vanishingly weak, testing the model’s sensitivity and its ability to avoid vanishing gradients.
@@
    Half-Cube Map. This distribution challenges models with heavy-tailed data. We begin with two correlated Gaussian variables and apply the homeomorphism $\text{half-cube}(z)=|z|^{3/2}\text{sign}(z)$ to each. The resulting variables, ${\mathbf{x}}^{\prime}=\text{half-cube}({\mathbf{x}})$ and ${\mathbf{y}}^{\prime}=\text{half-cube}({\mathbf{y}})$, have significantly heavier tails than the original Gaussians. While this transformation preserves the mutual information, i.e., $I({\mathbf{x}}^{\prime};{\mathbf{y}}^{\prime})=I({\mathbf{x}};{\mathbf{y}})$, the distorted, heavy-tailed geometry can be difficult for non-parametric methods that rely on local density estimation.
@@
    Asinh Mapping. This task tests the model’s ability to handle distributions with highly concentrated densities. We start with two independent standard Gaussian random variables, $\mathbf{x}$} and $\mathbf{y}$}, and apply the inverse hyperbolic sine transformation, $\text{asinh}(z)=\log(z+\sqrt{z^{2}+1})$, to each. The resulting variables, ${\mathbf{x}}^{\prime}=\text{asinh}({\mathbf{x}})$ and ${\mathbf{y}}^{\prime}=\text{asinh}({\mathbf{y}})$, form our two distributions. This mapping compresses the tails of the original Gaussian distributions into a central region with extremely high curvature, creating sharp density peaks that can cause gradient instability and pose significant challenges for kernel-based methods.
@@
    Additive Noise. This scenario is designed to evaluate performance on distributions with fragmented support and sharp discontinuities. Let $\mathbf{x}$} be a random variable uniformly distributed on $[0,1]$, i.e., ${\mathbf{x}}\sim\mathcal{U}(0,1)$, and let $\mathbf{n}$} be an independent noise variable, ${\mathbf{n}}\sim\mathcal{U}(-\epsilon,\epsilon)$. We define the second variable as ${\mathbf{y}}={\mathbf{x}}+{\mathbf{n}}$. The resulting joint distribution is piecewise-constant and possesses sharp, non-differentiable boundaries. This structure violates the smoothness assumptions underlying many score-based estimators, providing a stringent test of model robustness. The true MI is given by $I({\mathbf{x}};{\mathbf{y}})=\log(2\epsilon)+0.5$ for $\epsilon\leq 0.5$.
@@
    Gamma-Exponential. This task features a complex, non-linear dependency structure. We define a hierarchical relationship where the first variable $\mathbf{x}$} is drawn from a Gamma distribution, ${\mathbf{x}}\sim\text{Gamma}(\rho,1)$. The second variable $\mathbf{y}$} is then drawn from an Exponential distribution whose rate parameter is given by the value of $\mathbf{x}$}, i.e., ${\mathbf{y}}\mid{\mathbf{x}}=x\sim\text{Exponential}(x)$. The resulting joint distribution is highly asymmetric and non-Gaussian. The true mutual information is $I({\mathbf{x}};{\mathbf{y}})=\psi(\rho+1)-\log(\rho)$, where $\psi$ is the digamma function.
The results in Tab. 6(e) can be summarized as follows: (1) On the Edge-Singular Gaussian task, MinPV (both affine and spherical) attains the lowest or near-lowest MSE, especially under extreme correlations ($\rho\to 0,\pm 1$). On the Half-Cube Map task with heavy tails, the spherical constraint yields stable and accurate estimates, outperforming all baselines. (2) On the Asinh Mapping task with concentrated densities and sharp curvature, MinPV again achieves the lowest error, while baselines such as VP struggle with density peaks. (3) The largest gains appear on the Additive Noise and Gamma–Exponential tasks. With sharp discontinuities and complex non-linear dependencies, fixed-path methods degrade severely, whereas MinPV, particularly with the spherical constraint, achieves dramatically lower MSE, often by an order of magnitude.
Overall, no fixed path is uniformly effective across all five benchmarks, with performance varying by data geometry. In contrast, MinPV consistently delivers state-of-the-art accuracy by optimizing path variance, producing data-adaptive trajectories tailored to each problem’s structure.
Table 6: Full MSE results for MI estimation on five geometrically pathological datasets. The top row of each sub-table indicates the varying correlation coefficient $\rho$. Our MinPV path demonstrates consistently superior or competitive performance across the wide range of challenging data geometries, highlighting the benefits of a learnable path.
@@
MinPV (ours) affine $0.0010$ $\mathbf{0.0002}$ $0.0005$ $0.0005$ $\mathbf{0.0001}$ $\mathbf{0.0003}$ $\mathbf{0.0001}$ $0.0001$ $\mathbf{0.0000}$ $0.0002$ $\mathbf{0.0001}$ $\mathbf{0.0001}$ $\mathbf{0.0001}$ $\mathbf{0.0000}$ $0.0001$ $\mathbf{0.0001}$ $\mathbf{0.0002}$ $0.0007$ $\mathbf{0.0001}$
MinPV (ours) spherical $\mathbf{0.0005}$ $0.0006$ $\mathbf{0.0001}$ $0.0003$ $0.0004$ $0.0004$ $0.0002$ $\mathbf{0.0000}$ $0.0001$ $\mathbf{0.0001}$ $0.0002$ $\mathbf{0.0001}$ $0.0004$ $0.0001$ $\mathbf{0.0000}$ $0.0004$ $0.0005$ $0.0006$ $0.0007$
@@
MinPV (ours) affine $\mathbf{0.0005}$ $\mathbf{0.0004}$ $0.0006$ $\mathbf{0.0003}$ $\mathbf{0.0004}$ $0.0006$ $0.0007$ $0.0007$ $0.0008$ $0.0008$ $0.0026$ $0.0014$ $0.0025$ $\mathbf{0.0002}$ $0.0007$ $0.0004$ $0.0004$ $0.0004$ $0.0013$
MinPV (ours) spherical $0.0005$ $0.0005$ $\mathbf{0.0002}$ $0.0006$ $0.0008$ $\mathbf{0.0002}$ $\mathbf{0.0003}$ $\mathbf{0.0005}$ $\mathbf{0.0002}$ $0.0012$ $\mathbf{0.0002}$ $0.0004$ $\mathbf{0.0006}$ $0.0006$ $\mathbf{0.0002}$ $\mathbf{0.0003}$ $\mathbf{0.0003}$ $\mathbf{0.0003}$ $0.0014$
@@
MinPV (ours) affine $0.0003$ $\mathbf{0.0003}$ $0.0004$ $\mathbf{0.0002}$ $0.0001$ $\mathbf{0.0003}$ $\mathbf{0.0005}$ $\mathbf{0.0000}$ $0.0002$ $\mathbf{0.0004}$ $\mathbf{0.0001}$ $0.0003$ $0.0005$ $0.0005$ $0.0005$ $0.0005$ $\mathbf{0.0001}$ $0.0003$ $\mathbf{0.0004}$
MinPV (ours) spherical $\mathbf{0.0002}$ $\mathbf{0.0003}$ $\mathbf{0.0003}$ $0.0007$ $\mathbf{0.0001}$ $0.0007$ $0.0010$ $\mathbf{0.0000}$ $\mathbf{0.0001}$ $0.0009$ $0.0006$ $\mathbf{0.0000}$ $\mathbf{0.0001}$ $0.0007$ $0.0005$ $\mathbf{0.0003}$ $0.0004$ $\mathbf{0.0002}$ $0.0006$
@@
MinPV (ours) affine $0.2360$ $0.0199$ $0.0081$ $0.0167$ $0.0161$ $0.0216$ $0.0304$ $0.0155$ $\mathbf{0.0010}$
MinPV (ours) spherical $\mathbf{0.1016}$ $\mathbf{0.0134}$ $\mathbf{0.0078}$ $\mathbf{0.0009}$ $\mathbf{0.0138}$ $\mathbf{0.0117}$ $\mathbf{0.0196}$ $\mathbf{0.0026}$ $0.0029$
@@
MinPV (ours) affine $0.4785$ $1.5466$ $0.1137$ $0.1274$ $0.0915$ $\mathbf{0.0063}$ $\mathbf{0.0032}$ $0.0294$ $0.0069$
MinPV (ours) spherical $\mathbf{0.2721}$ $\mathbf{0.2304}$ $\mathbf{0.0917}$ $\mathbf{0.0434}$ $\mathbf{0.0075}$ $0.0187$ $0.0091$ $\mathbf{0.0030}$ $\mathbf{0.0004}$
@@ B.5 Experimental Settings and Results for Density Estimation
Figure 4: Density estimation on six structured and multi-modal datasets. MinPV successfully learns a data-adaptive path tailored to the specific manifold of each dataset.
@@ B.6 Illustration of Optimized Path Schedules
We visualize the path schedules learned under the MinPV principle across different tasks and datasets. Here, “prior” and “posterior” denote the initialized and optimized path schedules.

### v3

@@ 6 Conclusions and Future Works
This paper addresses the key paradox in score-based density ratio estimation (DRE) that estimator performance strongly depends on the chosen path schedule while being theoretically path-invariant. Our analysis reveals that this discrepancy arises from an overlooked path variance term in training objectives. Based on this observation, we established the Minimum Path Variance (MVP) principle, arguing that this term should be explicitly minimized rather than ignored. We then proposed MVP, a framework that derives a closed-form expression for path variance and directly optimizes it. By parameterizing the path with a Kumaraswamy Mixture (KMM) Model, MVP learns a data-adaptive low-variance path. This eliminates the need for heuristic path selection and ensures closer alignment with the ideal score-matching objective. Experiments demonstrate that MVP achieves SOTA performance in estimation error and stability across challenging DRE benchmarks.
@@
The MVP principle is general and could benefit other score-based generative models by learning optimal noise schedules to improve sample quality and stability. Beyond accuracy, future research could focus on computational efficiency, for example by leveraging optimal transport–based path formulations or adopting lighter parameterizations to further accelerate training. We believe that explicitly optimizing the geometry of interpolation paths will continue to afford significant advances in density ratio estimation and related score-based methods.
@@ A.1 Assumptions
A.5 Control of the Lipschitz Constant via Minimum Variance Path Principle
@@ B.1 Formulations and Variances of Common Path Schedules
This path is inspired by the variance-preserving (VP) SDE schedule (Ho et al., 2020; Song et al., 2021b), satisfying the spherical constraint $\alpha(t)^{2}+\beta(t)^{2}=1$. We use the standard diffusion constants $\beta_{0}=0.1$ and $\beta_{1}=20$.
@@ B.3 Experimental Settings for the Kumaraswamy Distribution
PyTorch implementation of MVP path with KMM parameterization.
We also provide the PyTorch implementation of the proposed MVP in this paper, as shown in Algorithm 3.
Algorithm 3 PyTorch implementation of MVP path with KMM parameterization.
1class MVPPath($n$.Module):
@@
3 An implementation for computing the MVP path schedules
@@ B.4 Experimental Settings and Results for $f$-divergence Estimation
    Edge-Singular Gaussian. This task assesses model stability in edge-case correlation scenarios. We consider two jointly Gaussian random variables, $({\mathbf{x}},{\mathbf{y}})\sim\mathcal{N}(\mathbf{0},[[1,\rho],[\rho,1]])$. The mutual information, given by $\text{MI}({\mathbf{x}};{\mathbf{y}})=-0.5\log(1-\rho^{2})$, becomes numerically challenging at the extremes of the correlation coefficient $\rho$. As $\rho\to 1^{-}$, the distributions become nearly degenerate and collinear, leading to numerical divergence in many estimators. Conversely, as $\rho\to 0^{+}$, the dependency signal becomes vanishingly weak, testing the model’s sensitivity and its ability to avoid vanishing gradients.
@@
    Half-Cube Map. This distribution challenges models with heavy-tailed data. We begin with two correlated Gaussian variables and apply the homeomorphism $\text{half-cube}({\mathbf{z}})=|{\mathbf{z}}|^{3/2}\mathrm{sign}({\mathbf{z}})$ to each. The resulting variables, ${\mathbf{x}}^{\prime}=\text{half-cube}({\mathbf{x}})$ and ${\mathbf{y}}^{\prime}=\text{half-cube}({\mathbf{y}})$, have significantly heavier tails than the original Gaussians. While this transformation preserves the mutual information, i.e., $\text{MI}({\mathbf{x}}^{\prime};{\mathbf{y}}^{\prime})=\text{MI}({\mathbf{x}};{\mathbf{y}})$, the distorted, heavy-tailed geometry can be difficult for non-parametric methods that rely on local density estimation.
@@
    Asinh Mapping. This task tests the model’s ability to handle distributions with highly concentrated densities. We start with two independent standard Gaussian random variables, $\mathbf{x}$} and $\mathbf{y}$}, and apply the inverse hyperbolic sine transformation, $\mathrm{asinh}({\mathbf{z}})=\log({\mathbf{z}}+\sqrt{{\mathbf{z}}^{2}+1})$, to each. The resulting variables, ${\mathbf{x}}^{\prime}=\mathrm{asinh}({\mathbf{x}})$ and ${\mathbf{y}}^{\prime}=\mathrm{asinh}({\mathbf{y}})$, form our two distributions. This mapping compresses the tails of the original Gaussian distributions into a central region with extremely high curvature, creating sharp density peaks that can cause gradient instability and pose significant challenges for kernel-based methods.
@@
    Additive Noise. This scenario is designed to evaluate performance on distributions with fragmented support and sharp discontinuities. Let $\mathbf{x}$} be a random variable uniformly distributed on $[0,1]$, i.e., ${\mathbf{x}}\sim\mathcal{U}({\bm{0}},{\bm{1}})$, and let $\mathbf{n}$} be an independent noise variable, ${\mathbf{n}}\sim\mathcal{U}(-\epsilon,\epsilon)$. We define the second variable as ${\mathbf{y}}={\mathbf{x}}+{\mathbf{n}}$. The resulting joint distribution is piecewise-constant and possesses sharp, non-differentiable boundaries. This structure violates the smoothness assumptions underlying many score-based estimators, providing a stringent test of model robustness. The true MI is given by $\text{MI}({\mathbf{x}};{\mathbf{y}})=\log(2\epsilon)+0.5$ for $\epsilon\leq 0.5$.
@@
    Gamma-Exponential. This task features a complex, non-linear dependency structure. We define a hierarchical relationship where the first variable $\mathbf{x}$} is drawn from a Gamma distribution, ${\mathbf{x}}\sim\mathrm{Gamma}(\rho,1)$. The second variable $\mathbf{y}$} is then drawn from an Exponential distribution whose rate parameter is given by the value of $\mathbf{x}$}, i.e., ${\mathbf{y}}\mid{\mathbf{x}}={\bm{x}}\sim\mathrm{Exponential}({\bm{x}})$. The resulting joint distribution is highly asymmetric and non-Gaussian. The true mutual information is $I({\mathbf{x}};{\mathbf{y}})=\psi(\rho+1)-\log(\rho)$, where $\psi$ is the digamma function.
The results in Tab. 6(e) can be summarized as follows: (1) On the Edge-Singular Gaussian task, MVP (both affine and spherical) attains the lowest or near-lowest MSE, especially under extreme correlations ($\rho\to 0,\pm 1$). On the Half-Cube Map task with heavy tails, the spherical constraint yields stable and accurate estimates, outperforming all baselines. (2) On the Asinh Mapping task with concentrated densities and sharp curvature, MVP again achieves the lowest error, while baselines such as VP struggle with density peaks. (3) The largest gains appear on the Additive Noise and Gamma–Exponential tasks. With sharp discontinuities and complex non-linear dependencies, fixed-path methods degrade severely, whereas MVP, particularly with the spherical constraint, achieves dramatically lower MSE, often by an order of magnitude.
Overall, no fixed path is uniformly effective across all five benchmarks, with performance varying by data geometry. In contrast, MVP consistently delivers state-of-the-art accuracy by optimizing path variance, producing data-adaptive trajectories tailored to each problem’s structure.
Table 6: Full MSE results for MI estimation on five geometrically pathological datasets. The top row of each sub-table indicates the varying correlation coefficient $\rho$. Our MVP path demonstrates consistently superior or competitive performance across the wide range of challenging data geometries, highlighting the benefits of a learnable path.
@@
MVP (ours) affine $0.0010$ $\mathbf{0.0002}$ $0.0005$ $0.0005$ $\mathbf{0.0001}$ $\mathbf{0.0003}$ $\mathbf{0.0001}$ $0.0001$ $\mathbf{0.0000}$ $0.0002$ $\mathbf{0.0001}$ $\mathbf{0.0001}$ $\mathbf{0.0001}$ $\mathbf{0.0000}$ $0.0001$ $\mathbf{0.0001}$ $\mathbf{0.0002}$ $0.0007$ $\mathbf{0.0001}$
MVP (ours) spherical $\mathbf{0.0005}$ $0.0006$ $\mathbf{0.0001}$ $0.0003$ $0.0004$ $0.0004$ $0.0002$ $\mathbf{0.0000}$ $0.0001$ $\mathbf{0.0001}$ $0.0002$ $\mathbf{0.0001}$ $0.0004$ $0.0001$ $\mathbf{0.0000}$ $0.0004$ $0.0005$ $0.0006$ $0.0007$
@@
MVP (ours) affine $\mathbf{0.0005}$ $\mathbf{0.0004}$ $0.0006$ $\mathbf{0.0003}$ $\mathbf{0.0004}$ $0.0006$ $0.0007$ $0.0007$ $0.0008$ $0.0008$ $0.0026$ $0.0014$ $0.0025$ $\mathbf{0.0002}$ $0.0007$ $0.0004$ $0.0004$ $0.0004$ $0.0013$
MVP (ours) spherical $0.0005$ $0.0005$ $\mathbf{0.0002}$ $0.0006$ $0.0008$ $\mathbf{0.0002}$ $\mathbf{0.0003}$ $\mathbf{0.0005}$ $\mathbf{0.0002}$ $0.0012$ $\mathbf{0.0002}$ $0.0004$ $\mathbf{0.0006}$ $0.0006$ $\mathbf{0.0002}$ $\mathbf{0.0003}$ $\mathbf{0.0003}$ $\mathbf{0.0003}$ $0.0014$
@@
MVP (ours) affine $0.0003$ $\mathbf{0.0003}$ $0.0004$ $\mathbf{0.0002}$ $0.0001$ $\mathbf{0.0003}$ $\mathbf{0.0005}$ $\mathbf{0.0000}$ $0.0002$ $\mathbf{0.0004}$ $\mathbf{0.0001}$ $0.0003$ $0.0005$ $0.0005$ $0.0005$ $0.0005$ $\mathbf{0.0001}$ $0.0003$ $\mathbf{0.0004}$
MVP (ours) spherical $\mathbf{0.0002}$ $\mathbf{0.0003}$ $\mathbf{0.0003}$ $0.0007$ $\mathbf{0.0001}$ $0.0007$ $0.0010$ $\mathbf{0.0000}$ $\mathbf{0.0001}$ $0.0009$ $0.0006$ $\mathbf{0.0000}$ $\mathbf{0.0001}$ $0.0007$ $0.0005$ $\mathbf{0.0003}$ $0.0004$ $\mathbf{0.0002}$ $0.0006$
@@
MVP (ours) affine $0.2360$ $0.0199$ $0.0081$ $0.0167$ $0.0161$ $0.0216$ $0.0304$ $0.0155$ $\mathbf{0.0010}$
MVP (ours) spherical $\mathbf{0.1016}$ $\mathbf{0.0134}$ $\mathbf{0.0078}$ $\mathbf{0.0009}$ $\mathbf{0.0138}$ $\mathbf{0.0117}$ $\mathbf{0.0196}$ $\mathbf{0.0026}$ $0.0029$
@@
MVP (ours) affine $0.4785$ $1.5466$ $0.1137$ $0.1274$ $0.0915$ $\mathbf{0.0063}$ $\mathbf{0.0032}$ $0.0294$ $0.0069$
MVP (ours) spherical $\mathbf{0.2721}$ $\mathbf{0.2304}$ $\mathbf{0.0917}$ $\mathbf{0.0434}$ $\mathbf{0.0075}$ $0.0187$ $0.0091$ $\mathbf{0.0030}$ $\mathbf{0.0004}$
@@ B.5 Experimental Settings and Results for Density Estimation
Figure 4: Density estimation on six structured and multi-modal datasets. MVP successfully learns a data-adaptive path tailored to the specific manifold of each dataset.
@@ B.6 Illustration of Optimized Path Schedules
We visualize the path schedules learned under the MVP principle across different tasks and datasets. Here, “prior” and “posterior” denote the initialized and optimized path schedules.